OpenStudy (anonymous):

In EFG, mE = 30, mF = 60, and mG = 90. Which of the following statements about EFG are true?

OpenStudy (anonymous):

Check all that apply. A. EG = sqrt 3 * FG B. EG = sqrt 3 /2 * EF C. FG = sqrt 3/2 EF D. EF = 2 * FG E. EG = 2 * FG F. EF = sqrt 3 FG

OpenStudy (anonymous):

@mathstudent55

OpenStudy (raden):